# Transpiling quantum circuits by a transformers-based algorithm

### v1

@@ I Introduction
A transformer capable of accurately translating quantum circuits between IBM and IonQ gate sets, achieving over $99.98\%$ accuracy with scalable complexity, is developed. Quantum circuits are ordered sequences of logic gates acting on qubits. The most conventional way to implement such gates is via unitary operators, even though another feasible approach, in quantum mechanics, is through disruptive measurements (the so called one-way, or measurement-based quantum computing) [corli2025quantum, browne2016one, barenco1995elementary]. Since current devices operate in the noisy intermediate-scale quantum (NISQ) regime, the practical execution of a circuit depends crucially on its efficiency. Fewer qubits and shorter gate sequences reduce the accumulation of hardware noise and increase the fidelity of the computation. The process of transforming a circuit into a form that is both efficient and physically executable is known as quantum compiling [maronese2022quantum, khaneja2001cartan, corli2025gauge]. Quantum compiling involves transformations that preserve the logical behavior of a quantum algorithm while improving or adapting its circuit representation. Such process can reduce the depth of the circuit, break down complex operations, or restructure the circuit to meet the physical constraints of a specific hardware platform. In fact, each quantum technology (superconducting qubits [arute2019quantum], trapped ions [haffner2008quantum, blatt2012quantum], neutral atoms [henriet2020quantum], spin in semiconductors [de2023silicon] or photonic systems [o2007optical, takeda2017universal]) offers a distinct set of native gates determined by the underlying physical interactions that it can directly implement. Therefore, a circuit written for one platform generally can not be executed on another without being adapted and re-expressed. Such translation, known as quantum transpiling [hua2023qasmtrans, kamaka2020quantum, wesley2024linguaquanta], converts a circuit from one set of gates to an equivalent sequence compatible with the target hardware and its native gates. As in natural language translation, the structure of the sequence can change significantly, but its meaning, i.e., the quantum algorithm encoded by the logic circuit, must remain exactly the same.
@@
Use of transformer models in quantum computing workflows has been recently proposed [russo2025optimizing, ruiz2025quantum]. Early applications include their use in variational quantum eigensolvers (VQE), where transformers help identify ground state configurations of parametrized quantum circuits [nakaji2024generative]. They have also been employed to study and predict the expressiveness of quantum ansatze, providing insight into how circuit architectures explore the underlying Hilbert space [zhang2025learning]. Beyond heuristic or variational approaches, models such as Ket-GPT have demonstrated the ability to generate complete sequences of quantum logic gates, illustrating the potential of large language models to operate directly on circuit-level representations [apak2024ketgpt]. Some of us has previously studied data-driven strategies for quantum compiling through complementary approaches, including evolutionary techniques based on genetic algorithms [de2015universal], reinforcement learning frameworks capable of autonomously optimizing circuit structures [moro2021quantum, semola2022deep], and gauge theory methods designed to apply structural constraints and improve compilation efficiency [corli2024measurement, corli2025gauge2]. Here, instead, we focus on the problem of translating circuits between the native gate sets of different hardware platforms using transformer architecture. In this work, we demonstrate that a transformer model is capable of accurately translating quantum circuits of up to five qubits between distinct vocabularies of quantum gates, achieving an accuracy rate that consistently exceeds 99.98%. The model learns to convert circuits expressed in the native gate set of the IBM hardware to the native gate set used by the trapped-ion quantum processors of IonQ. To enable the translation, we represent each circuit using OpenQASM, the hardware-independent quantum assembly language supported by Qiskit and widely adopted as the textual standard for describing quantum circuits. The transformer operates directly on this textual representation, taking the OpenQASM description of a source circuit and generating the corresponding sequence of gates compatible with the target device. The model effectively performs circuit-level translation in a manner analogous to natural language translation, but within the structured symbolic domain of quantum gate operations.
@@ II Implementation of the transformer-based transpiler
Transformer architecture has recently become the leading model for sequential data processing in machine learning. Using stacked encoder and decoder blocks, transformer architecture maps an input sequence to a high-dimensional latent representation and generates the corresponding output sequence [lin2022survey]. Through self-attention mechanisms, transformers capture long-range dependencies by projecting tokens into a vector space where semantic and structural relationships are encoded. The encoder identifies the relevant features of the input, while the decoder uses such information to produce a coherent and contextually appropriate output.
@@ II.1.2 Tuning the tokenizer for the QASM standard
Although often treated as a secondary component in architecture design, the tokenizer is the key building block of this work. Indeed, deep learning models cannot process raw text, since they require numerical inputs. The tokenizer bridges this gap by breaking text down into atomic units, called tokens, and converting them into numerical representations which the model can encode and learn from. In standard Natural Language Processing (NLP), a balance must be struck [vaswani2017attention]. Assigning a token to every distinct word creates an unmanageably large vocabulary, while assigning a token to every letter strips away any semantic meaning. Therefore, NLP usually finds an optimal middle ground using sub-word units.
@@ II.1.3 QASM Embedding
After the tokenization step, which converts the QASM code into sequences of integer indices, the data is ready to be fed into the model. The initial stage of the architecture – and a key factor in the success of Transformers in NLP – is the projection of these discrete tokens into a learnable latent space. Such latent space is the internal geometric representation of the model in which raw inputs (sequences of text for Transformers or images for convolutional neural networks [li2021survey], graphs or nodes for graph neural networks [zhou2020graph, lee2019mathematical] and so on) are mapped to vectors where each component is a real value. These continuous vectors capture the essential structure and meaning of the inputs by simply assigning a position in a huge dimensional space. In fact, similar entities are mapped and located close to each other, allowing the model to reason about relationships, patterns, and context through Euclidean geometry, i.e., distances and directions, as shown in the second step of Figure 1. During training, backpropagation refines this mapping to produce token embeddings. These dense vectors serve as the foundational input fed into the Transformer layers.
@@ II.2 Training loss and accuracy setting
To prevent the model from becoming overconfident and to encourage better generalization, we apply label smoothing [muller2019does] to the cross-entropy objective. Instead of training on hard one-hot targets (where the correct token probability is exactly 1.0), we smooth the target distribution using a factor $\epsilon$. The new target probability $y^{LS}_{c}$ for the correct class becomes $1-\epsilon$, while the remaining probability mass $\epsilon$ is distributed uniformly across the vocabulary:
@@ II.2.1 Metrics to assess the correctness of the training
When dealing with NLP, the perplexity is a known metric defined as the exponential expression for the cross-entropy $\mathcal{L}_{CE}$ defined in Equation (4) [jurafsky2020speech, pg.38]:
@@
Such metric quantifies the statistical surprisal [modirshanechi2022taxonomy] of the model regarding the target data. Intuitively, a perplexity value of $k$ indicates that the model is, on average, as uncertain as if it were choosing uniformly among $k$ candidates for the next token. A perfect training would entail $\hat{y}_{i}=1\,\forall i$, therefore the perplexity would be equal to $1$. Consequently, a lower perplexity signifies that the model is assigning higher probability mass to the correct tokens. The goal during training is to drive this value as close as possible to the theoretical minimum of 1, which would represent absolute certainty and perfect prediction.
@@ II.3 Benchmarking parametric continuous circuits versus a universal discrete set of gates
with $d(\cdot,\cdot)$ being a proper distance in the space of the unitary matrices $SU(2)$, e.g. the trace distance or the diamond norm [nielsen2010quantum]. The total number of tokens, required to characterize a quantum logic gate, can be instead described as
@@ Appendix D The Solovay-Kitaev scaling laws for the tokenization process
The Solovay-Kitaev theorem states that, once a discrete subset $\mathcal{G}\subset SU(2)$ of logic gates, dense in $SU(2)$, is given, any single-qubit unitary $U$ can be approximated with precision $\epsilon$ in $O(\log^{c}(1/\epsilon))$ gates from $\mathcal{G}$ [nielsen2010quantum, dawson2005solovay, maronese2022quantum]. The constant $c$ is assessed to range between $2$ and $4$, depending on the formulation of the theorem [dawson2005solovay]. This theorem plays a crucial role in quantum compiling theory, since it states that, in order to implement a fault-tolerant gate $U(\theta)$ through a set of native gates, it takes only a polylogarithmic number of gates, avoiding worse scaling (such as a polynomial or even an exponential one). The Solovay-Kitaev theorem allows to convert any circuit in a discrete set of fault-tolerant gates – for instance, the Clifford gates and the $S$ one – at the cost of introducing a certain degree of precision $\epsilon$ when approximating the single-qubit rotations [dawson2005solovay]. Moreover, the theorem do not need to address a specific universal set of gates, since any of universal set fits the hypothesis of the theorem. Such a statement means that, once a specific hardware display its own universal set of native logic gates, the Solovay-Kitaev theorem can be applied via such a suitable set [maronese2022quantum]. In other words, the Solovay-Kitaev theorem is agnostic about the hardware and the platform we are considering.

### v2

@@ I Introduction
A transformer capable of accurately translating quantum circuits between IBM and IonQ gate sets, achieving over $99.98\%$ accuracy with scalable complexity, is developed. Quantum circuits are ordered sequences of logic gates acting on qubits. The most conventional way to implement such gates is via unitary operators, even though another feasible approach, in quantum mechanics, is through disruptive measurements (the so called one-way, or measurement-based quantum computing) [6, 5, 3]. Since current devices operate in the noisy intermediate-scale quantum (NISQ) regime, the practical execution of a circuit depends crucially on its efficiency. Fewer qubits and shorter gate sequences reduce the accumulation of hardware noise and increase the fidelity of the computation. The process of transforming a circuit into a form that is both efficient and physically executable is known as quantum compiling [22, 18, 8]. Quantum compiling involves transformations that preserve the logical behavior of a quantum algorithm while improving or adapting its circuit representation. Such process can reduce the depth of the circuit, break down complex operations, or restructure the circuit to meet the physical constraints of a specific hardware platform. In fact, each quantum technology (superconducting qubits [2], trapped ions [13, 4], neutral atoms [14], spin in semiconductors [12] or photonic systems [28, 32]) offers a distinct set of native gates determined by the underlying physical interactions that it can directly implement. Therefore, a circuit written for one platform generally can not be executed on another without being adapted and re-expressed. Such translation, known as quantum transpiling [15, 17, 34], converts a circuit from one set of gates to an equivalent sequence compatible with the target hardware and its native gates. As in natural language translation, the structure of the sequence can change significantly, but its meaning, i.e., the quantum algorithm encoded by the logic circuit, must remain exactly the same.
@@
Use of transformer models in quantum computing workflows has been recently proposed [30, 29]. Early applications include their use in variational quantum eigensolvers (VQE), where transformers help identify ground state configurations of parametrized quantum circuits [26]. They have also been employed to study and predict the expressiveness of quantum ansatze, providing insight into how circuit architectures explore the underlying Hilbert space [35]. Beyond heuristic or variational approaches, models such as Ket-GPT have demonstrated the ability to generate complete sequences of quantum logic gates, illustrating the potential of large language models to operate directly on circuit-level representations [1]. Some of us has previously studied data-driven strategies for quantum compiling through complementary approaches, including evolutionary techniques based on genetic algorithms [11], reinforcement learning frameworks capable of autonomously optimizing circuit structures [24, 31], and gauge theory methods designed to apply structural constraints and improve compilation efficiency [7, 9]. Here, instead, we focus on the problem of translating circuits between the native gate sets of different hardware platforms using transformer architecture. In this work, we demonstrate that a transformer model is capable of accurately translating quantum circuits of up to five qubits between distinct vocabularies of quantum gates, achieving an accuracy rate that consistently exceeds 99.98%. The model learns to convert circuits expressed in the native gate set of the IBM hardware to the native gate set used by the trapped-ion quantum processors of IonQ. To enable the translation, we represent each circuit using OpenQASM, the hardware-independent quantum assembly language supported by Qiskit and widely adopted as the textual standard for describing quantum circuits. The transformer operates directly on this textual representation, taking the OpenQASM description of a source circuit and generating the corresponding sequence of gates compatible with the target device. The model effectively performs circuit-level translation in a manner analogous to natural language translation, but within the structured symbolic domain of quantum gate operations.
@@ II Implementation of the transformer-based transpiler
Transformer architecture has recently become the leading model for sequential data processing in machine learning. Using stacked encoder and decoder blocks, transformer architecture maps an input sequence to a high-dimensional latent representation and generates the corresponding output sequence [21]. Through self-attention mechanisms, transformers capture long-range dependencies by projecting tokens into a vector space where semantic and structural relationships are encoded. The encoder identifies the relevant features of the input, while the decoder uses such information to produce a coherent and contextually appropriate output.
@@ II.1.2 Tuning the tokenizer for the QASM standard
Although often treated as a secondary component in architecture design, the tokenizer is the key building block of this work. Indeed, deep learning models cannot process raw text, since they require numerical inputs. The tokenizer bridges this gap by breaking text down into atomic units, called tokens, and converting them into numerical representations which the model can encode and learn from. In standard Natural Language Processing (NLP), a balance must be struck [33]. Assigning a token to every distinct word creates an unmanageably large vocabulary, while assigning a token to every letter strips away any semantic meaning. Therefore, NLP usually finds an optimal middle ground using sub-word units.
@@ II.1.3 QASM Embedding
After the tokenization step, which converts the QASM code into sequences of integer indices, the data is ready to be fed into the model. The initial stage of the architecture – and a key factor in the success of Transformers in NLP – is the projection of these discrete tokens into a learnable latent space. Such latent space is the internal geometric representation of the model in which raw inputs (sequences of text for Transformers or images for convolutional neural networks [20], graphs or nodes for graph neural networks [36, 19] and so on) are mapped to vectors where each component is a real value. These continuous vectors capture the essential structure and meaning of the inputs by simply assigning a position in a huge dimensional space. In fact, similar entities are mapped and located close to each other, allowing the model to reason about relationships, patterns, and context through Euclidean geometry, i.e., distances and directions, as shown in the second step of Figure 1. During training, backpropagation refines this mapping to produce token embeddings. These dense vectors serve as the foundational input fed into the Transformer layers.
@@ II.2 Training loss and accuracy setting
To prevent the model from becoming overconfident and to encourage better generalization, we apply label smoothing [25] to the cross-entropy objective. Instead of training on hard one-hot targets (where the correct token probability is exactly 1.0), we smooth the target distribution using a factor $\epsilon$. The new target probability $y^{LS}_{c}$ for the correct class becomes $1-\epsilon$, while the remaining probability mass $\epsilon$ is distributed uniformly across the vocabulary:
@@ II.2.1 Metrics to assess the correctness of the training
When dealing with NLP, the perplexity is a known metric defined as the exponential expression for the cross-entropy $\mathcal{L}_{CE}$ defined in Equation (4) [16, pg.38]:
@@
Such metric quantifies the statistical surprisal [23] of the model regarding the target data. Intuitively, a perplexity value of $k$ indicates that the model is, on average, as uncertain as if it were choosing uniformly among $k$ candidates for the next token. A perfect training would entail $\hat{y}_{i}=1\,\forall i$, therefore the perplexity would be equal to $1$. Consequently, a lower perplexity signifies that the model is assigning higher probability mass to the correct tokens. The goal during training is to drive this value as close as possible to the theoretical minimum of 1, which would represent absolute certainty and perfect prediction.
@@ II.3 Benchmarking parametric continuous circuits versus a universal discrete set of gates
with $d(\cdot,\cdot)$ being a proper distance in the space of the unitary matrices $SU(2)$, e.g. the trace distance or the diamond norm [27]. The total number of tokens, required to characterize a quantum logic gate, can be instead described as
@@ Appendix D The Solovay-Kitaev scaling laws for the tokenization process
The Solovay-Kitaev theorem states that, once a discrete subset $\mathcal{G}\subset SU(2)$ of logic gates, dense in $SU(2)$, is given, any single-qubit unitary $U$ can be approximated with precision $\epsilon$ in $O(\log^{c}(1/\epsilon))$ gates from $\mathcal{G}$ [27, 10, 22]. The constant $c$ is assessed to range between $2$ and $4$, depending on the formulation of the theorem [10]. This theorem plays a crucial role in quantum compiling theory, since it states that, in order to implement a fault-tolerant gate $U(\theta)$ through a set of native gates, it takes only a polylogarithmic number of gates, avoiding worse scaling (such as a polynomial or even an exponential one). The Solovay-Kitaev theorem allows to convert any circuit in a discrete set of fault-tolerant gates – for instance, the Clifford gates and the $S$ one – at the cost of introducing a certain degree of precision $\epsilon$ when approximating the single-qubit rotations [10]. Moreover, the theorem do not need to address a specific universal set of gates, since any of universal set fits the hypothesis of the theorem. Such a statement means that, once a specific hardware display its own universal set of native logic gates, the Solovay-Kitaev theorem can be applied via such a suitable set [22]. In other words, the Solovay-Kitaev theorem is agnostic about the hardware and the platform we are considering.